Problem 3. n chromium(II) oxide, CrO, is obtained from pure elements, the composition of solid is actually Cro9201.00, and the compound is electrically neutral, and dont conduct electricity. Provide an explanation for this fact.
Answer #1

The compound is electrically neutral. It means total number of positive charges is equal to total number of negative charges. The charge on O ions is -2. Some Cr ions have charge of +2. Some Cr ions have charge of +3. For 100 oxide ions, there are 76 Cr(II) ions and 16 Cr(III) ions. Total 92 chromium ions are present for 100 oxide ions. This gives formula Cr_{0.92}O_{1.00} .

Total charge on 100 oxide ions =-2 \times 100 = -200

Total charge on 76 Cr(II) ions =+2 \times 76 =+152

Total charge on 16 Cr(III) ions =+3 \times 16 =+48

Total positive charges =+152+48=+200

Thus total charges cations is equal to total charges on anions. This maintains electrical neutrality.